I really need help on this problem.
The directions say for each function-list the range.
1. K:T------> 12/T+1 DOMAIN = {1,2,3}
2. F:Z------> 3(Z-1)^2-8 DOMAIN = {-1,0,1,2,3}
3. H(X) = X^2-3X+1 DOMAIN = {-1,0,1}
I have no clue how to do these. Could you do at least one for me. Thank you.

Answer by stanbon(75887)

I really need help on this problem.
The directions say for each function-list the range.
1. K:T------> 12/T+1 DOMAIN = {1,2,3}
Solve using each of the domain values:
1---------------12/1 + 1 = 12+1 = 13
2---------------12/2 + 1 = 6 + 1 = 7
3---------------12/3 + 1 = 4 + 1 = 5
Then the range is {13,7,5}
======================================================
2. F:Z------> 3(Z-1)^2-8 DOMAIN = {-1,0,1,2,3}
-1------------3(-1-1)^2-8 = 3(-2)^2-8 = 3*4- 8 = 4
0-------------3(0-1)^2-8 = 3(1)-8 = 2-8 = -5
1-------------3(1-1)^2-8 = 3*0-8 = -8
2-------------3(2-1)^2-8 = 3*1-8 = -5
3-------------3(3-1)^2-8 = 3*4-8 = 4
Then the range is {4,-5,-8,}
======================================================
Hope that helps.
Cheers,
Stan H.
3. H(X) = X^2-3X+1 DOMAIN = {-1,0,1}
I have no clue how to do these. Could you do at least one for me. Thank you.